Write the thermochemical equation for the following reactions that occur at constant temperature:



a) Decomposition of mercuric oxide, HgO, to mercury and oxygen gas, requires 87.1 kJ of heat




b) Combustion of hexane, C6H12, to carbon dioxide gas and water, releases 3923.9 kJ of heat

1
Expert's answer
2022-04-27T06:37:04-0400

a. 2HgO = 2Hg + O2 - 87.1 kJ ∆H = 87.1 kJ

b. C6H12 + 9O2 = 6CO2 + 6H2O + 3923.9 kJ ∆H = 3923.9 kJ